LHC 2010 - latest news

Week 29: Machine coordinators: Oliver Bruening and Ralph Assmann

Main aims of the week:

Latest plan: link PDF link XLS (version of 21 June, 21:30)

List of End of Fill studies (in work)

List of MD requests (in work)

Day Summary Link to report from morning meeting
Monday 26th
  • 00h50: Start ramp, 1b per beam, 9e10 per bunch. Heavy losses. B1 lost in ramp.
  • 01h38: Beam 2 at 3.5 TeV: loss map tests.
  • 02h08: Requested beam dump.
  • 03h00: Injection.
  • 04h36: Start ramp, 1b per beam, 9e10 per bunch. Heavy losses. B1 lost in ramp.
  • 05h19: RCS.A45B1 tripped again twice (second time at 06h07). Need access for repair.
  • 08h00: Access for RCS.A45B1 repair and repair of temperature sensor on RQD.A23.
Slides from 8h30 meeting
Sunday 25th
  • 00h34: Beam dump test at 3.5 TeV with pilot bunch.
  • 02h15: Injection studies, interupted by various HW problems. Beta beat measurement after new 10 A/s cycling.
  • 07h00: Reset of P2 compensator (problem during night) and pre-cycle.
  • 09h30: Prepare for test ramp with pilot bunch (fix chromaticity and orbit problems).
  • 10h26: Start of ramp.
  • 11h23: Requested beam dump at 3.5 TeV. Rampdown and precycle. CMS cosmic run. Ramp up of ALICE solenoid. Work on BLM crate in IR7.
  • 13h00: Injection. Orbit feedback problem (uncontrolled trims).
  • 15h12: Start ramp, 1b per beam, 9e10 per bunch. Heavy losses.
  • 16h12: SMP tests completed, including 3.5 TeV check. Requested dump.
  • 17h38: Work on BLM crate in IR7: temporary measures (autmoatic reboot) implemented.
  • 17h39: Trip RCS.A45B1. Access for repair. Temperature sensor on RQTL9.R7B2. Pre-cycle.
  • 23h15: Injection setup.
Slides from 9h00 meeting
Saturday 24th
  • 04h00: SPS magnet intervention completed. CMS access.
  • 08h00: PS RF problem.
  • 12h30: Beam back at LHC. Injection setup for pilot bunch.
  • 14h23: Beam dump check - IR6 BPM interlock
  • 17h05: Injection check for nominal bunch.
  • 18h20: Repair attempt for BLM crate "cfv-sr7-blmr".
  • 23h12: Prepare ramp.
  • 23h41: Test ramp for beam dump (pilot bunch).
Slides from 9h00 meeting
Friday 23rd
  • 06h30: Start of LHC dry runs with 10 A/s.
  • 08h00: End of SPS MD's.
  • 11h00: Short in SPS main dipole diagnosed. Need to replace main dipole in SPS. NO BEAM for LHC. Continuing LHC dry runs.
Slides from 8h30 meeting
Slides from LHC Status Report (LPPC)
Thursday 22nd
  • 18h00: LHC machine closed. End of technical stop for LHC.
  • LHC recovery. Requalification of various magnet circuits (following interventions during technical stop).
  • No beam from SPS: MD's until 08h00 Friday morning.


No slides during technical stop
Wednesday 21st

Technical stop

No slides during technical stop
Slides from LMC report

Tuesday 20th

Technical stop (news Katy Foraz):

  • Activities are progressing according to schedule, and even ahead of schedule: Non conformities on cryo valves are fixed, the "electrical un-lock-out phase 2" is done at points 1, 3 and 5
  • Problems which occured
    • Leaks in UA63 - inside a power converter - normal situation will be back on wednesday morning
    • PM45 lift unavailable during morning,
    • Patrols lost !!
No slides during technical stop
Monday 19th

Technical stop (news Katy Foraz):

  • The RP surveys are completed. The LHC machine tunnel is classified as Supervised Radiation Area. Exceptions are LSS3, LSS7 and the UD caverns at Point 6 which are classified as Simple Controlled Area with local postings at the most radioactive elements. No surveys have been performed in TI2/8.
  • Electrical lock-out are done
  • All the activities are progressing according to schedule
  • Problems:
    • Unforeseen lost of patrols: arc 45, arc 34, RUX45, UX45, arc 78 due to a fault of the access system !!!
    • Problem with server – unavailability of access system
Report from 8:30 meeting